OpenStudy (anonymous):

solve the equation. 6(w-5)-3w=w+10

OpenStudy (anonymous):

distribute the 6 then combine like terms

OpenStudy (anonymous):

i'm confused lol

OpenStudy (anonymous):

just look at 6(w-5) to start. Can you distribute this portion of the equation?

OpenStudy (anonymous):

6*w=6w-5 ?

OpenStudy (anonymous):

or is it 6w-30?

OpenStudy (anonymous):

correct. so now what would you do?

OpenStudy (anonymous):

wait which one? lol

OpenStudy (anonymous):

is i

OpenStudy (anonymous):

is it 6w-30?

OpenStudy (anonymous):

yes but that is only part of the equation\[6w-30-3w=w+10\] can you solve from here by combining like terms?

OpenStudy (anonymous):

subract 6w from the 6w and 3w?